# Joint Beamforming and Antenna Position Optimization for Fluid Antenna-Assisted MU-MIMO Networks

### v1

@@ V-B Impact of Box-Constrained Movement Mode
We evaluate the impact of the box-constrained movement mode on system performance. The parameter $\rho$ is used to control the size of the movable region for each FA. Specifically, when $\rho=0.5$, each FA is restricted to movement along the $y$-axis for the box-constrained movement mode. As $\rho$ increases, the movable region expands, allowing greater flexibility.
Fig. 3 presents the simulation results, which is averaged over $10^{3}$ samples with $M=16$. The WSR performance gap between systems with and without the box-constrained movement mode decreases as $\rho$ increases. Notably, when $\rho\geq 2$, the performance gap remains below $0.3\%$. This indicates that when the movable region of each FA is sufficiently large, the impact of the box-constrained movement mode on performance is negligible. This is because a larger movable region for each FA increases the likelihood of avoiding deep fading channels. Based on this observation, we only test the performance with the box-constrained movement mode, and set $\rho=2$ for all subsequent simulations.
V-C Convergence Behavior
Figure 4: Convergence behavior of the proposed algorithms.
Fig. 4 illustrates the convergence behavior of our proposed algorithms. Regardless of the number of transmit FAs $M$, all algorithms converge within $80$ iterations. Moreover, all curves increase with the number of iterations, confirming the convergence of our proposed algorithms.
The decentralized implementation experiences a slight performance loss compared with the centralized approach, especially in the initial iterations. However, the total number of iterations required for convergence remains similar. Moreover, the performance gap is less than $1\%$ after convergence, making the difference negligible. Therefore, we evaluate only the decentralized implementation for all subsequent simulations.
V-D Performance Comparison
TABLE I: WSR Performance Comparison
@@ V-D Performance Comparison
FPA RPA TFA RFA TRFA (ours)
$16$ $20$ $0.864$ $0.830$ $1.30$ $1.11$ $\mathbf{1.55}$
$30$ $3.30$ $3.28$ $4.15$ $3.89$ $\mathbf{4.62}$
$64$ $20$ $2.12$ $2.07$ $2.78$ $2.38$ $\mathbf{3.11}$
$30$ $6.68$ $6.81$ $7.69$ $7.26$ $\mathbf{8.27}$
$256$ $20$ $4.30$ $4.31$ $5.18$ $4.67$ $\mathbf{5.62}$
$30$ $12.8$ $13.2$ $13.7$ $13.6$ $\mathbf{14.5}$
We compare the WSR of TRFA optimized by the decentralized implementation of our proposed algorithm with several baselines under different configurations, as shown in Table I. The performance of TRFA is consistently the highest across all configurations, demonstrating its superior performance. Specifically, a larger transmit FA number $M$ and a higher power budget $P_{\max}$ lead to a significant absolute WSR increase. Meanwhile, when the transmit FA number $M$ is smaller and the power budget $P_{\max}$ is lower, the relative WSR improvement over FPA is more pronounced. With $M=64$ and $P_{\max}=20$ dBm, the WSR increases by approximately $47\%$. For $M=256$ and $P_{\max}=30$ dBm, the WSR gain over FPA is about $1.7$ bps/Hz.
The effectiveness of our proposed optimization algorithm is further validated by comparing the WSR values of TRFA and RPA. The WSR of RPA is similar to that of FPA, indicating that random FA position adjustments hardly provide any performance gain. In contrast, TRFA achieves significantly higher WSR than FPA, demonstrating that our algorithm effectively optimizes FA positions to enhance performance.
V-E Impact of Power Budget and Number of Users
@@
Fig. 5 illustrates the WSR performance with different transmit power budgets, where we set $M=64$. The WSR of all systems increases significantly with a higher transmit power budget, and TRFA consistently outperforms the baselines. This is because the FA can dynamically reconstruct the channel, improving the receive SINR with the same transmit power. Therefore, utilizing FA can significantly reduce the required transmit power. Specifically, by fixing the WSR at $2$ bps/Hz, the transmit power budget can be reduced by over $3$ dB, demonstrating the superior performance of the FA-assisted MU-MIMO system.
Meanwhile, the impact of the number of users is shown in Fig. 6 for $M=64$. The WSR of TRFA consistently exceeds that of all baselines, demonstrating its superior performance in multiuser systems. More importantly, as the number of users $K$ increases, the performance gap between TRFA and the baselines widens. With more users, conventional MU-MIMO systems with FPAs experience stronger MUI. In contrast, FAs can be dynamically repositioned to mitigate MUI, thereby enhancing system capacity.
V-F Computational Efficiency of Decentralized Implementation
TABLE II: CPU Time Saved Compared with Centralized
@@ V-F Computational Efficiency of Decentralized Implementation
$M$ $C$
FPA
TFA
RFA
TRFA
$64$ $4$ $89.6\%$ $64.5\%$ $89.3\%$ $69.4\%$
$16$ $98.3\%$ $89.4\%$ $90.4\%$ $87.9\%$
$256$ $4$ $99.2\%$ $68.6\%$ $97.8\%$ $71.9\%$
$16$ $99.5\%$ $80.5\%$ $97.8\%$ $82.4\%$
$\frac{\partial f_{\rm Quad}}{\partial p_{m}^{t}}=\sum_{k=1}^{K}\mathtt{tr}\left(\left(\frac{\partial f_{\rm Quad}}{\partial\mathbf{G}_{k}}\right)^{\mathsf{T}}\frac{\partial\mathbf{G}_{k}}{\partial p_{m}^{t}}+\frac{\partial\mathbf{G}_{k}^{\mathsf{H}}}{\partial p_{m}^{t}}\left(\frac{\partial f_{\rm Quad}}{\partial\mathbf{G}_{k}^{\mathsf{H}}}\right)^{\mathsf{T}}\right)=2\sum_{k=1}^{K}\Re\left\{\mathtt{tr}\left(\left(\frac{\partial f_{\rm Quad}}{\partial\mathbf{G}_{k}}\right)^{\mathsf{T}}\frac{\partial\mathbf{G}_{k}}{\partial p_{m}^{t}}\right)\right\}$ (81)
$[\frac{\partial\mathbf{g}_{k}\left(\mathbf{t}_{m}\right)}{\partial x_{m}^{t}}]_{q}$ $=\jmath\frac{2\pi}{\lambda}\cos\theta_{kq}^{t}\cos\phi_{kq}^{t}\exp\left[\jmath\frac{2\pi}{\lambda}\left(x_{m}^{t}\cos\theta_{kq}^{t}\cos\phi_{kq}^{t}+y_{m}^{t}\cos\theta_{kq}^{t}\sin\phi_{kq}^{t}+z_{m}^{t}\sin\theta_{kq}^{t}\right)\right],$ (83a)
$[\frac{\partial\mathbf{g}_{k}\left(\mathbf{t}_{m}\right)}{\partial y_{m}^{t}}]_{q}$ $=\jmath\frac{2\pi}{\lambda}\cos\theta_{kq}^{t}\sin\phi_{kq}^{t}\exp\left[\jmath\frac{2\pi}{\lambda}\left(x_{m}^{t}\cos\theta_{kq}^{t}\cos\phi_{kq}^{t}+y_{m}^{t}\cos\theta_{kq}^{t}\sin\phi_{kq}^{t}+z_{m}^{t}\sin\theta_{kq}^{t}\right)\right],$ (83b)
$[\frac{\partial\mathbf{g}_{k}\left(\mathbf{t}_{m}\right)}{\partial z_{m}^{t}}]_{q}$ $=\jmath\frac{2\pi}{\lambda}\sin\theta_{kq}^{t}\exp\left[\jmath\frac{2\pi}{\lambda}\left(x_{m}^{t}\cos\theta_{kq}^{t}\cos\phi_{kq}^{t}+y_{m}^{t}\cos\theta_{kq}^{t}\sin\phi_{kq}^{t}+z_{m}^{t}\sin\theta_{kq}^{t}\right)\right].$ (83c)
$\bigg{\lvert}\frac{\partial^{2}f_{\rm Quad}}{\partial p_{m}^{t}\partial p_{j}^{t}}\bigg{\rvert}$ $=2\Bigg{\lvert}\sum_{k=1}^{K}\Re\left\{-[\hat{\mathbf{W}}]_{mj}\left(\frac{\partial\mathbf{g}_{k}^{\mathsf{H}}\left(\mathbf{t}_{j}\right)}{\partial p_{j}^{t}}\right)^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\left(\frac{\partial\mathbf{g}_{k}\left(\mathbf{t}_{m}\right)}{\partial p_{m}^{t}}\right)+\delta_{mj}[\mathbf{D}_{k}^{\mathtt{Tx}}]_{m}\frac{\partial^{2}\mathbf{g}_{k}\left(\mathbf{t}_{m}\right)}{\partial p_{m}^{t}\partial p_{j}^{t}}\right\}\Bigg{\rvert}$ (84a)
$\leq\frac{8\pi^{2}}{\lambda^{2}}\sum_{k=1}^{K}L_{k}^{t}\left(\left(\Big{\lvert}[\hat{\mathbf{W}}]_{mj}\Big{\rvert}+\sqrt{M}\delta_{mj}\Big{\lVert}[\hat{\mathbf{W}}]_{m}\Big{\rVert}_{2}\right)\right.$
×$\hskip 45.00008pt\left.\times\lVert\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\rVert_{2}+\delta_{mj}\sqrt{\frac{\alpha_{k}}{L_{k}^{t}}}\Big{\lVert}[\overline{\mathbf{W}}_{k}]_{m}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big{\rVert}_{2}\right)\triangleq\gamma^{\mathtt{Tx}}_{mj}.$ (84b)
We assess the computational efficiency of the proposed decentralized implementation by measuring the central processing unit (CPU) time of each algorithm. Specifically, for centralized implementations, the total CPU time is recorded. For decentralized implementations, the CPU time is calculated as the sum of the CU’s CPU time and the maximum running time among all DUs.
The results in Table II show that decentralized implementations significantly reduce computation time compared with centralized implementations across all system configurations. The proposed decentralized approach demonstrates high computational efficiency, achieving at least a $69.4\%$ reduction in CPU time. This improvement is attributed to the parallel processing capability of DUs, where each DU independently solves a smaller-scale problem. As the number of DUs $C$ increases, the problem size of each DU decreases, leading to greater parallelism. Consequently, for a fixed number of transmit FAs $M$, the CPU time reduction becomes more pronounced. However, increasing the number of DUs also introduces higher hardware complexity and interconnection costs, highlighting a tradeoff between computational efficiency and system complexity.
VI Conclusion
In this paper, we considered the joint beamforming and antenna position optimization for WSR maximization in FA-assisted MU-MIMO systems. We decoupled the original WSR maximization problem using matrix FP techniques and employed BCA to solve the resulting subproblems. To optimize FA positions, we proposed a novel MM algorithm capable of optimizing all antenna positions simultaneously. To further reduce computational, storage, and interconnection costs, we proposed a decentralized implementation of our algorithm based on the DBP architecture. Simulation results demonstrated that FA-assisted MU-MIMO networks optimized by our algorithms achieve significant performance gains across various setups. Moreover, the decentralized implementation substantially improves computational efficiency.
-A Derivation of $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$ and $\delta^{\mathtt{Tx}}$
@@ VI Conclusion
Using the chain rule, we compute the entries of $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$ via (81) at the bottom of the page, where the calculation of $\frac{\partial f_{\rm Quad}}{\partial\mathbf{G}_{k}}$ and $\frac{\partial\mathbf{G}_{k}}{\partial p_{m}^{t}}$ are given by (29) and
@@
$\frac{\partial\mathbf{G}_{k}}{\partial p_{m}^{t}}=\left[\mathbf{0},\cdots,\mathbf{0},\frac{\partial\mathbf{g}_{k}\left(\mathbf{t}_{m}\right)}{\partial p_{m}^{t}},\mathbf{0},\cdots,\mathbf{0}\right],$ (82)
@@
respectively. Here, the $q$-th entries of $\frac{\partial\mathbf{g}_{k}\left(\mathbf{t}_{m}\right)}{\partial p_{m}^{t}},\ p\in\{x,y,z\}$, is given by (83) at the bottom of the page. Plugging (29) and (82) into (81), we obtain the expression of $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$.
We utilize the inequality (26) to derive the closed-form expression for $\delta^{\mathtt{Tx}}$. Specifically, we approximate the maximum eigenvalue of the Hessian matrix $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}^{2}f_{\rm Quad}(\mathbf{T})$ using an upper bound on its matrix infinity norm [40]. To compute this upper bound, we first determine the maximum absolute value of each entry in the Hessian matrix. The absolute values of these entries are given by (84a) at the top of the next page. Next, we derive an upper bound on $\big{\lVert}[\mathbf{D}_{k}^{\mathtt{Tx}}]\big{\rVert}_{2}$ for any $\mathbf{T}$, which provides an upper bound for (84a). The resulting upper bound on $\big{\lVert}[\mathbf{D}_{k}^{\mathtt{Tx}}]\big{\rVert}_{2}$ is given by (85) at the top of the next page.
$\Big{\lVert}[\mathbf{D}_{k}^{\mathtt{Tx}}]_{m}\Big{\rVert}_{2}\leq\sqrt{\alpha_{k}}\Big{\lVert}[\overline{\mathbf{W}}_{k}]_{m}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}^{\mathsf{H}}\Big{\rVert}_{2}+\sqrt{ML_{k}^{t}}\Big{\lVert}[\hat{\mathbf{W}}]_{m}\Big{\rVert}_{2}\Big{\lVert}\mathbf{\Sigma}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}\overline{\mathbf{\Phi}}_{k}\left(\mathbf{I}+\overline{\mathbf{\Gamma}}_{k}\right)\overline{\mathbf{\Phi}}_{k}^{\mathsf{H}}\overline{\mathbf{F}}_{k}^{\mathsf{H}}\mathbf{\Sigma}_{k}\Big{\rVert}_{\rm F}.$ (85)
By substituting (85) into (84a), we derive the upper bounds of the absolute values of the entries in $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}^{2}f_{\rm Quad}(\mathbf{T})$, given by (84b) at the top of the next page. For simplicity, we define the upper bounds of $\big{\lvert}\frac{\partial^{2}f_{\rm Quad}}{\partial p_{m}^{t}\partial p_{j}^{t}}\big{\rvert}$ as $\gamma^{\mathtt{Tx}}_{mj}$. Finally, the upper bound of $\lambda_{\max}\left(\nabla^{2}_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}(\mathbf{T})\right)$ is obtained by computing the upper bound of $\big{\lVert}\nabla^{2}_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}(\mathbf{T})\big{\rVert}_{\infty}$, given by

### v2

@@ VI-A Impact of Box-Constrained Movement Mode
We evaluate the impact of the box-constrained movement mode on system performance. The parameter $\rho$ is used to control the size of the movable region for each FA. Specifically, when $\rho=0.5$, each FA is restricted to movement along the $y$-axis for the box-constrained movement mode. As $\rho$ increases, the movable region expands, allowing greater flexibility. Fig. 6 presents the simulation results with $M=16$. The WSR performance gap between systems with and without the box-constrained movement mode decreases as $\rho$ increases. Notably, when $\rho\geq 2$, the performance gap remains below $1\%$. This indicates that once the movable region of each FA is sufficiently large, the impact of the box-constrained movement mode on performance is negligible. This is because a larger $\rho$ provides sufficient spatial DoF under both movement modes, resulting in comparable capabilities to avoid deep fading channels. Based on this observation, we only test the performance with the box-constrained movement mode, and set $\rho=2$ for all subsequent simulations.
VI-B Convergence Behavior
Figure 7: Convergence behaviors of the proposed algorithms.
Fig. 7 illustrates the convergence behaviors of our proposed algorithms. Regardless of the number of transmit FAs $M$, all algorithms converge within $80$ iterations. It is also observed that the decentralized implementation converges slightly slower than the centralized approach, especially in the initial iterations. Nonetheless, the performance gap becomes negligible after convergence.
VI-C Performance Comparison
1) Comparison with Prior Art [31]
Figure 8: WSR comparison with [31] under MU-MISO.
We compare the WSR performance of the proposed centralized and decentralized algorithms with the conventional method in [31]. Since the method in [31] is designed for MU-MISO systems, we simplify our system by setting $K=1$ and $d=1$. The simulation results are presented in Fig. 8. For FPA, the WMMSE algorithm in [31] and the FP-based algorithm proposed in this paper achieve nearly identical performance, which is expected due to the equivalence between WMMSE and FP [27, 28]. For TRFA, both the proposed centralized and decentralized algorithms outperform the conventional method. The WSR gain is attributed to the proposed parallel MM algorithm, which constructs tighter surrogate functions and enables more effective optimization.
2) Comparison with Baselines
TABLE II: WSR Performance Comparison
@@ VI-C Performance Comparison
C-FPA D-FPA C-RPA D-RPA C-TFA D-TFA C-RFA D-RFA C-TRFA D-TRFA
$16$ $30$ $0.682$ $0.682$ $0.640$ $0.641$ $1.10$ $1.10$ $0.908$ $0.909$ $1.33$ $1.34$
$40$ $3.02$ $3.03$ $2.98$ $2.98$ $4.01$ $3.99$ $3.64$ $3.64$ $4.52$ $4.51$
$64$ $30$ $1.76$ $1.76$ $1.69$ $1.69$ $2.51$ $2.51$ $2.04$ $2.05$ $2.87$ $2.87$
$40$ $6.53$ $6.53$ $6.58$ $6.58$ $7.83$ $7.74$ $7.15$ $7.15$ $8.47$ $8.38$
$256$ $30$ $4.06$ $4.06$ $4.00$ $4.00$ $5.10$ $5.07$ $4.40$ $4.41$ $5.58$ $5.55$
$40$ $13.2$ $13.2$ $13.6$ $13.6$ $14.4$ $14.2$ $14.0$ $14.0$ $15.3$ $15.0$
We compare the WSR performance of TRFA, optimized by both the proposed centralized and decentralized implementations, against several baselines under various configurations, as summarized in Table II. The performance of C-TRFA and D-TRFA is consistently the highest across all configurations, demonstrating that the proposed BCA-based algorithm can exploit the spatial DoF of the system. The effectiveness of the proposed MM algorithm is validated by comparing the WSR values of TRFA and RPA. The WSR of RPA is similar to that of FPA, indicating that random FA position adjustments hardly provide any performance gain. In contrast, TRFA achieves significantly higher WSR than FPA, demonstrating that the proposed MM algorithm effectively optimizes FA positions to enhance performance. Compared with the centralized implementation, the decentralized implementation achieves a similar performance. For FPA, the WSR of the decentralized algorithm is no worse than that of the centralized algorithm. For TFA, the maximum WSR loss of the decentralized implementation is $1.41\%$ when $M=256$ and $P_{\max}=40$ dBm. For TRFA, the maximum WSR loss of the decentralized implementation is $2.00\%$ when $M=256$ and $P_{\max}=40$ dBm as well.
3) Impact of Power Budget and Number of Users
@@
Fig. 9 illustrates the WSR performance with different transmit power budgets. The WSR of all systems increases significantly with a higher transmit power budget, and TRFA consistently outperforms the baselines. The WSR of the decentralized implementation is similar to that of their centralized counterparts. The improved WSR performance of TRFA is attributed to the ability of FAs to dynamically reconstruct the channel, thereby enhancing the receive SINR under a fixed transmit power budget. As a result, FAs can significantly reduce the required transmit power to achieve a target performance. By fixing the WSR at $2$ bps/Hz, the transmit power budget can be reduced by around $4$ dB, demonstrating the superior performance of the FA-assisted MU-MIMO system.
The impact of the number of users is shown in Fig. 10. The WSR performance of TRFA consistently exceeds that of all baselines, and the decentralized implementation achieves a similar WSR as the centralized implementation. More importantly, as the number of users $K$ increases, the performance gap between TRFA and the baseline methods becomes more pronounced. In conventional MU-MIMO systems with FPAs, a larger number of users leads to more severe MUI, thereby degrading system performance. In contrast, the dynamically repositioning of FAs allows effective MUI mitigation, which in turn enhances the overall system capacity.
VI-D Robust Analysis
Throughout the paper, we assume that CSI is perfectly known at the BS. This assumption, however, may not hold in practice due to channel estimation errors. In this part, we evaluate the effect of CSI errors on the system performance. First, we evaluate the impact of AoA/AoD errors on WSR. Denote the estimated elevation and azimuth AoD as $\tilde{\theta}_{kq}^{\mathtt{Tx}}$ and $\tilde{\phi}_{kq}^{\mathtt{Tx}}$, respectively, and the estimated elevation and azimuth AoA as $\tilde{\theta}_{kq}^{\mathtt{Rx}}$ and $\tilde{\phi}_{kq}^{\mathtt{Rx}}$, respectively. The difference between the estimated AoA/AoD and the ground truth AoA/AoD follows a uniform distribution, i.e., $\tilde{\theta}_{kq}^{\mathtt{Tx}}-\theta_{kq}^{\mathtt{Tx}}\sim\mathcal{U}\left[-\mu_{\theta,\phi},\mu_{\theta,\phi}\right]$, $\tilde{\phi}_{kq}^{\mathtt{Tx}}-\phi_{kq}^{\mathtt{Tx}}\sim\mathcal{U}\left[-\mu_{\theta,\phi},\mu_{\theta,\phi}\right]$, $\tilde{\theta}_{kq}^{\mathtt{Rx}}-\theta_{kq}^{\mathtt{Rx}}\sim\mathcal{U}\left[-\mu_{\theta,\phi},\mu_{\theta,\phi}\right]$, and $\tilde{\phi}_{kq}^{\mathtt{Rx}}-\phi_{kq}^{\mathtt{Rx}}\sim\mathcal{U}\left[-\mu_{\theta,\phi},\mu_{\theta,\phi}\right]$, where $\mu_{\theta,\phi}$ is the maximum AoA/AoD error. The simulation result, shown in Fig. 11, demonstrates that TRFA and TFA are more sensitive to the AoA/AoD errors than FPA and RPA. The reason is that inaccurate AoA/AoD may mislead the FAs to position themselves on undesirable channels, leading to performance degradation.
Figure 11: WSR versus the AoA/AoD error $\mu_{\theta,\phi}$.
Figure 12: WSR versus the PRM error $\varepsilon_{\mathbf{\Sigma}}$.
Then, we evaluate the impact of PRM errors on WSR. We represent the estimated PRM as $\tilde{\mathbf{\Sigma}}_{k}$. The normalized difference of each entry between the estimated PRM and the ground truth PRM follows a CSCG distribution, i.e., $\frac{[\tilde{\mathbf{\Sigma}}_{k}]_{qq}-[\mathbf{\Sigma}_{k}]_{qq}}{\lvert[\mathbf{\Sigma}_{k}]_{qq}\rvert}\sim\mathcal{CN}\left(0,\varepsilon_{\mathbf{\Sigma}}\right)$, where $\varepsilon_{\mathbf{\Sigma}}$ is the PRM error. As shown in Fig. 12, all schemes are similarly robust to the PRM error. Even with very high PRM errors, i.e., $\varepsilon_{\mathbf{\Sigma}}=1$, the WSR of TRFA is still the highest among all schemes. The phenomenon consolidates the robustness of the proposed algorithm.
Remark: The proposed algorithm exhibits higher sensitivity to AoA/AoD errors compared to PRM errors. This is because small-scale fading, which FAS primarily exploits to enhance WSR, is more significantly influenced by AoA/AoD parameters. The simulation results further indicate that the proposed algorithm has a higher demand on the accuracy of AoA/AoD estimation than on PRM estimation.
VI-E Computational Efficiency of Decentralized Implementation
TABLE III: CPU Time Saved Compared with Centralized
@@ VI-E Computational Efficiency of Decentralized Implementation
$M$ $C$ FPA TFA RFA TRFA
$64$ 44 $95.5\%$ $67.3\%$ $77.8\%$ $65.6\%$
$16$ $96.1\%$ $77.1\%$ $78.4\%$ $73.1\%$
$256$ 44 $98.8\%$ $87.6\%$ $95.0\%$ $87.0\%$
$16$ $99.1\%$ $89.4\%$ $95.2\%$ $88.6\%$
We quantify the computational efficiency of the proposed algorithm by measuring central processing unit (CPU) time. Specifically, the total CPU time is recorded for the centralized algorithm, while for the decentralized algorithm under the DBP architecture, the CPU time is computed as the sum of the CU’s CPU time and the maximum running time among all DUs. As shown in Table III, the DBP architecture significantly reduces computation time by at least $65.6\%$ compared with the centralized algorithm across all system configurations. This improvement stems from the parallel processing capability of the DBP architecture, where each DU independently solves a smaller-scale problem. For a fixed number of transmit FAs $M$, increasing the number of DUs $C$ enhances parallelism, thereby saving more computational time.
VII Conclusion
In this paper, we investigated the joint beamforming and antenna position optimization problem for WSR maximization in FA-assisted MU-MIMO networks. To tackle the inherent coupling between beamforming matrices and antenna positions, we employed matrix FP techniques to decouple the problem and adopted the BCA framework to solve the resulting subproblems. For antenna position optimization, we proposed a novel parallel MM algorithm that enables simultaneous updates of all FA positions. To further reduce computational overhead, we developed a decentralized implementation based on the DBP architecture. Simulation results demonstrate that the proposed parallel MM algorithm significantly outperforms existing FA position optimization methods in terms of WSR performance. Moreover, FA-assisted MU-MIMO networks optimized by our algorithms achieve significant WSR gains across various setups compared with conventional MU-MIMO systems. The decentralized implementation achieves substantial reductions in computation time while maintaining performance that is nearly identical to its centralized counterpart. Additionally, we analyzed the robustness of the proposed algorithm to different types of channel uncertainty. The results reveal that the algorithm is more sensitive to AoA/AoD errors than to PRM errors, underscoring the importance of accurate angle estimation in FA-assisted MU-MIMO networks.
-A Derivation of $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$
@@ VII Conclusion
The entries of $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$ are computed using the matrix chain rule:
@@
$\frac{\partial f_{\rm Quad}}{\partial p_{m}^{\mathtt{Tx}}}$ $=\sum_{k=1}^{K}\mathtt{tr}\left[\left(\frac{\partial f_{\rm Quad}}{\partial\mathbf{G}_{k}}\right)^{\mathsf{T}}\frac{\partial\mathbf{G}_{k}}{\partial p_{m}^{\mathtt{Tx}}}+\frac{\partial\mathbf{G}_{k}^{\mathsf{H}}}{\partial p_{m}^{\mathtt{Tx}}}\left(\frac{\partial f_{\rm Quad}}{\partial\mathbf{G}_{k}^{\mathsf{H}}}\right)^{\mathsf{T}}\right]$
$=2\sum_{k=1}^{K}\Re\left\{\mathtt{tr}\left(\mathbf{D}_{k}^{\mathtt{Tx}}\frac{\partial\mathbf{G}_{k}}{\partial p_{m}^{\mathtt{Tx}}}\right)\right\},$ (77)
where $p\in\{x,y,z\}$ and $\mathbf{D}_{k}^{\mathtt{Tx}}$ is already given in (IV-D). Therefore, to compute $\frac{\partial f_{\rm Quad}}{\partial p_{m}^{\mathtt{Tx}}}$, it suffices to derive $\frac{\partial\mathbf{G}_{k}}{\partial p_{m}^{\mathtt{Tx}}}$:
$\frac{\partial\mathbf{G}_{k}}{\partial p_{m}^{\mathtt{Tx}}}=\Big[\underbrace{\mathbf{0},\dots,\mathbf{0}}_{m-1},\frac{\partial\mathbf{g}_{k}\left(\mathbf{t}_{m}\right)}{\partial p_{m}^{\mathtt{Tx}}},\underbrace{\mathbf{0},\dots,\mathbf{0}}_{M-m}\Big]\in\mathbb{C}^{L_{k}^{\mathtt{Tx}}\times M}.$ (78)
The $q$-th element of the partial derivative $\frac{\partial\mathbf{g}_{k}(\mathbf{t}_{m})}{\partial p_{m}^{\mathtt{Tx}}}$ can be compactly expressed as
$\left[\frac{\partial\mathbf{g}_{k}(\mathbf{t}_{m})}{\partial\mathbf{t}_{m}^{\mathsf{T}}}\right]_{q}$ $=\left[\left[\frac{\partial\mathbf{g}_{k}(\mathbf{t}_{m})}{\partial x_{m}^{\mathtt{Tx}}}\right]_{q},\left[\frac{\partial\mathbf{g}_{k}(\mathbf{t}_{m})}{\partial y_{m}^{\mathtt{Tx}}}\right]_{q},\left[\frac{\partial\mathbf{g}_{k}(\mathbf{t}_{m})}{\partial z_{m}^{\mathtt{Tx}}}\right]_{q}\right]$
$=\jmath\frac{2\pi}{\lambda}\left(\mathbf{g}_{kq}^{\mathtt{Tx}}\right)^{\mathsf{T}}\exp\left(\jmath\frac{2\pi}{\lambda}\rho_{kq}^{\mathtt{Tx}}(\mathbf{t}_{m})\right).$ (79)
@@
By substituting (-A) into (78), and then combining (IV-D) and (78) into (-A), we obtain the final expression of $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}\left(\mathbf{T}\right)$ given in (IV-D).
-B Derivation of $\delta^{\mathtt{Tx}}$
As indicated in (25), the constant $\delta^{\mathtt{Tx}}$ is constructed such that its value the upper bound of the maximum eigenvalue of Hessian matrix $\nabla_{\mathrm{vec}\left(\mathbf{T}\right)}^{2}f_{\rm Quad}(\mathbf{T})$. Since the calculation of eigenvalue is computationally expensive, we continue to find an upper bound of the maximum eigenvalue of the Hessian matrix to derive the closed-form expression for $\delta^{\mathtt{Tx}}$:
$\lambda_{\max}\left(\nabla^{2}_{\mathrm{vec}\left(\mathbf{T}\right)}f_{\rm Quad}(\mathbf{T})\right)$